Question. Given a ball with mass 0.6 kilograms hanging from a string with length 7 meters, what is the period of a pendulum swing when the swing doesn’t deviate more than π/12 radians from a free hang (where gravitational constant 9.8 meters per second)?
Answer #1

The time period will be given by

T = 2\pi\sqrt{\frac{l}{g}} = 2\times 3.14\times \sqrt{\frac{7}{9.8}} = 5.31s